OpenStudy (anonymous):

I really need help! Part 1. Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses. f(x)= x+a b g(x)=cx−d Part 2. Show your work to prove that the inverse of f(x) is g(x). Part 3. Show your work to evaluate g(f(x)). Part 4. Graph your two functions on a coordinate plane. Include a table of values for each function. Include 5 values for each function. Graph the line y = x on the same graph.

OpenStudy (danjs):

the inverse of a function, if there is one, flips the thing or mirrors it over the slant line y=x so for each point you can let y=x

OpenStudy (danjs):

for example for y= f(x) = x +a*b change the y and x, solving that will give you the reflection, inverse function f^(-1) x = y + a*b y = x - a*b inverse f^(-1)(x) = x - a*b

OpenStudy (anonymous):

so i put in what ever number like f(x)=5+2/3 and do i repeat the same numbers in g(x)=cx−d

OpenStudy (danjs):

oh i see, f(x) = (x+a)/b

OpenStudy (danjs):

inverse of that, change x and y around and resolve that for y x = (y+a)/b b*x = y + a y = b*x - a f^(-1)(x) that is the inverse f^(-1)(x) = bx - a

OpenStudy (danjs):

g(x) is cx - d g(x)=cx - d looks to be the same form, ig b=c and a=d

OpenStudy (anonymous):

thank i think i get it

OpenStudy (anonymous):

thank u not i sorry